OpenStudy (anonymous):

area addition postulate: find the area1 and area2 if x=3

OpenStudy (anonymous):

area addition postulate: find the area1 and area2 if x=3

OpenStudy (anonymous):

21*8+15*24 or not?

OpenStudy (anonymous):

where did you get that?

OpenStudy (anonymous):

i just put the 3 into x (from your attachment) and then you get your sides and multiply the sides. but it looks to easy for me too xD

OpenStudy (anonymous):

then what is the area1 and 2?

OpenStudy (anonymous):

area one is 21*8 and area two is 15*24

OpenStudy (anonymous):

how...

OpenStudy (anonymous):

Area (1) = (2+2x)(6x+3) = (2+2*3)(6*3+3) = 168 Area (2) = (7x+3)(4x+3) = (7*3+3)(4*3+3) = 360 Total area = Area (1) + Area (2) = 528

OpenStudy (anonymous):

thanks much!! @Zekarias